OpenStudy (anonymous):

Please Help give medal Factor: 4x^2-4x+1

OpenStudy (hoblos):

a^2 -2ab + b^2 = (a-b)^2 4x^2-4x+1 = (2x)^2 - 2(2x)(1) +1 = (2x-1)^2

OpenStudy (anonymous):

thank you.